OpenStudy (anonymous):

What is two-fifths divided by six-sevenths? twelve thirty-fifths two and eleven-twelfths seven-fifteenths two and one-seventh

OpenStudy (aaronq):

\(\dfrac{\dfrac{2}{5}}{\dfrac{6}{7}}=\dfrac{2*7}{6*5}=\dfrac{14}{30}=\dfrac{7}{15}\)

OpenStudy (anonymous):

thx

OpenStudy (aaronq):

np